# Scaling-aware rating of count forecasts

### v4

@@ 2.2 Corrections to the Poisson distribution
In practice, the Poisson assumption laid out above is never perfectly fulfilled. Several effects both increase and decrease the width of the distribution that one could ideally achieve to predict.
In the first place, perfectly knowing the number of visitors $N$ and the average probability $p$ to buy an item is a strong assumption – not all factors that influence demand can be known in practice. Hence, this item-level buy probability $N\cdot p$ is itself randomly distributed. If that distribution is a gamma-distribution, the resulting demand distribution becomes gamma-Poisson, that is, negative-binomial.
The forecast could, however, also exploit personal information about individual customers, that is, instead of assuming an average probability $p$ for the individual buy, this might be personalized to $p_{1}\dots p_{N}$. Le Cam’s theorem (Le Cam, 1960) sets a bound to the total variation distance between the resulting Poisson-binomial and Poisson distributions. Under plausible assumptions, the potential of added forecast sharpness via personalization is severely limited: An extraordinary degree of individualization (predictions of individual customer buy probabilities need to convey a lot of information, i.e. $p_{j}\ll 1$ or $p_{k}\approx 1$, for the same product) would be necessary to achieve a clearly-sharper-than-Poisson distribution.
The assumption that only one item is bought per customer – underlying the Bernoulli process – is also slightly broken in practice: At Sainsbury’s UK, about 87.5% of sold items are bought alone, 9.8% are bought in pairs of two pieces, about 1.5% in triplets of three pieces, four or more pieces are also occurring. The variance of the resulting distribution is thereby increased. In certain areas of retail, a buy of many pieces of the same item is not the exception but the rule: In hardware stores, bathroom tiles of a certain kind are not sold at all on most days, until someone wants to refurbish their bathroom and buys a large number. For such products, the Poisson assumption is clearly unsuitable. In grocery retail, however, the effect amounts to a manageable correction.
Finite stocks censor demand values that are larger than the available stocks. This right censoring effectively decreases the width of the observed distribution, which can give a false impression when evaluating a forecast for unconstrained demand against actual constrained sales.
A priori, it is difficult to estimate the strength of these effects. For the datasets that we employed – M5 and Sainsbury’s – we will see, ex post, that the Poisson limit remains a meaningful first-order benchmark, and that the effects mentioned in this section can be considered “second-order terms”, well behind precision scaling. Quantifying their impact more precisely is certainly desirable, and we plan to do so in future works. In general, if a more suitable benchmark distribution is available, which may include effects such as over-dispersion, zero-inflation, and right censoring, our method can easily be adapted, mutatis mutandis.
Hence, to focus on our main contribution, the scaling-aware rating of forecasts across selling rates, our working assumption is that a Poisson forecast constitutes the best possible, that is, the sharpest possible, prediction.
@@ 2.3 Forecasting model
Under the Poisson assumption, a product-location-day combination $j$ ($1\leq j\leq n$) is governed by a true selling rate $t_{j}$, which will forever remain unknown and is only under control in the numerical experiments conducted here. That is, the sales $s_{j}$ are random samples from the Poisson distribution with rate $\mu$ assuming the true value $t_{j}$:
@@
The forecast for that product-location-day $j$ is an estimated selling rate $r_{j}$, that is, the forecaster believes that each observed sales value $s_{j}$ behaves as if it were drawn from a Poisson distribution with rate $\mu=r_{j}$. The forecaster is self-confident and believes that the forecasted rates are indeed perfectly known Poisson-rates. Our considerations are fully independent of how the forecaster constructed the $r_{j}$: They might use a time-series approach, a machine-learning model, hybrid approaches, or any other method (Petropoulos, F. et al., 2022).
@@ 2.4 Evaluation metrics
To evaluate the quality of their forecast, the forecaster relates the observed sales $\vec{s}=(s_{1},\dots,s_{n})$ to their previously predicted rates $\vec{r}=(r_{1},\dots,r_{n})$ by evaluating a metric $M(\vec{r},\vec{s})$, where $M$ can take many different forms.
@@ 2.4.1 Loss functions on point estimates
The forecaster can use several loss functions to assess the forecast, including Mean Absolute Error (MAE),
@@
where the optimal point estimator for MAE, the median of the predicted distribution, replaces the mean $r_{j}$ (Schwertman et al., 1990; Gneiting, 2011). That is, the forecaster is aware that their forecast $r_{j}$ is not a universal best guess that can be fed into any metric, but that, depending on the chosen metric, an appropriate point estimate needs to be extracted (Kolassa, 2020). MAE naturally takes larger values for larger predictions, a first attempt towards a scale-free metric is the dimensionless normalized MAE (Kolassa and Schütz, 2007), often called Weighted Mean Absolute Percentage Error (WMAPE),
@@
The still-popular Mean Absolute Percentage Error (MAPE) comes with many flaws (Hyndman and Koehler, 2006; Kolassa and Martin, 2011; Kolassa and Schütz, 2007). In particular, it is undefined for observations $s=0$, whose treatment heavily affects the achieved MAPE-values (Kolassa, 2023a). These problems are complemented by the particularly complicated precision scaling of MAPE under the Poisson distribution described by Tichy (2023a), in which a simple removal of events with $s=0$ is assumed.
@@ 2.4.2 Ranked Probability Score
MAE has an immediate business interpretation: A planner who orders $\text{median}(r_{j})$ pieces will, on average, have an excess or under-stock of MAE pieces. The necessity to use the median of the distribution, however, leads to unpleasant discontinuous behavior (the median of the Poisson distribution jumps from 0 to 1 at the rate $r=\log 2\approx 0.693$). To provide a business interpretation while acknowledging the probabilistic nature of the forecast, the forecaster also evaluates the Mean Ranked Probability Score (MRPS), which is the mean over all product-location-days of the discrete Ranked Probability Score (RPS) (Epstein, 1969; Gneiting and Raftery, 2007; Tichy, 2022).
@@
MRPS generalizes MAE from simple point estimates to distributions: For a deterministic zero-width forecast that states that $r_{j}$ occurs with certainty, the first term in Eq. (7) becomes $|r_{j}-s_{j}|$ and second term vanishes (Gneiting and Raftery, 2007). We define a normalized version of MRPS, the Normalized Mean Ranked Probability Score (NMRPS),
@@
Precision scaling affects (N)MRPS in a similar way as MAE/WMAPE. The zoo of metrics is large, the sample that we chose here shall illustrate the ubiquity of precision scaling in different metrics and simultaneously cover the most popular ones.
2.5 Goals of forecast judgement
To make our discussion more tangible, we use the dataset for the validation period, 2016-04-25 to 2016-05-22, of the M5-competition dataset (Makridakis et al., 2022b, a). Our forecaster is interested in the overall performance of their model (“is that a good model?”), and in systematic patterns in the performance (“are we doing better in Hobbies or in Household?”), and therefore evaluates the forecast not only globally, but also segregated by department (or by other properties known at the moment of the forecast). Their goal is to judge the performance across strata to take high-level decisions: Shall data scientists focus on a certain product group, on a certain location, on certain patterns in time (Bower, 2023)? Which improvement can be expected? We aim at answering these questions in a non-defensive way that allows a clear operational interpretation (“the forecast is close/far from the Poisson ideal”), without falling into the naïve scaling trap, i.e., misinterpreting inevitable precision scaling effects due to the distribution’s heteroscedasticity as true differences in forecast quality.
@@ 2.6 Baseline model and naïve scaling trap
To obtain a maximally sharp and calibrated baseline model, we construct predictions $r^{\text{baseline}}_{j}$ via in-sample Expectation Maximization (EM) of the Poisson likelihood, applied on the entire set of sales $\vec{s}$ (see A). That is, the EM procedure ingests all outcomes $s_{j}$; using the empirical frequencies $P(s)$ and the Poisson distribution assumption (4), it computes a prior distribution of rates $P(r)$, from which it samples the rate $r^{\text{baseline}}_{j}$ for each outcome $s_{j}$ via $P(r|s)=P(s|r)P(r)/P(s)$. The rates $r_{j}$ that are then generated could hence have led to the outcome $s_{j}$ under a Poisson distribution. By construction, we thereby obtain a “perfect” model: For every set of rates around some value $r$, the associated observations seem to be sampled from a Poisson distribution of that rate $r$.
The Expectation Maximization procedure does not use any feature or any structure in the data (e.g., whether two sales values $s_{j}$, $s_{k}$ belong to the same product in different stores or to the same product-store-combination for different days is not accounted for). For our purpose of setting a benchmark, this baseline “forecast” (given that it is generated and evaluated using the same dataset, it is not, technically speaking, a forecast of some future event) $\vec{r}^{\text{baseline}}$ is suitable, since the sales $s_{j}$ are Poisson distributed for each prediction $r^{\text{baseline}}_{j}$.
As an alternative to using true sales data and artificially generated sales rates, one could have worked on a fully synthetic dataset. In that case, the selling rates themselves would be sampled randomly, e.g. from a gamma-distribution, and the observations $s_{j}$ would be Poisson-sampled for each rate $r_{j}$. Our point is, however, to exemplify the value of our method upon real-world datasets, which is why we use the M5 dataset as a starting point and need to generate an “ideal forecast”.
The calibration diagrams of Figure 1 confirm that the baseline $\vec{r}^{\text{baseline}}$ is an excellent model, since the unconditioned and the conditioned distributions of actually observed sales $s$ match the Poisson prediction well. It is fair to say that this baseline forecast is the best that one could possibly expect under the given set of product-related covariates (Tichy, 2023c). We will see in Section 5 that the state-of-the-art in commercial machine learning indeed comes close to the baseline performance.
Figure 1: Calibration diagrams for ideal baseline model, based on the heuristic of A: Distribution of observed sales frequency (orange) vs. predicted probability (blue, this color refers to the ideal Poisson case throughout this article). Upper left: Overall marginal calibration diagram (for visual convenience, we mapped the observation 0 to 0.5 to fit into the log-log-scale). Upper right: Cut on $2.21.8<r<$ 2.2. Lower left: Cut on $5.24.8<r<$ 5.2. Lower right: Cut on $10.29.8<r<$ 10.2.
@@
In Table 1, for the total, and for each of three selected departments (“Household 1”, “Hobbies 2”, “Foods 3”), all achieved metrics are finite when evaluated on the baseline forecast. Remember, however, that the forecast was fabricated to be ideally Poissonian for the entire assortment – by construction, it has the same quality in every subset of data. We know beforehand the result of a meaningful analysis: It should produce the result “there is no significant difference in forecast quality between the departments”.
Hence, the following intuitive conclusions are erroneous: “We need to improve Foods 3! Fast-movers are performing worse than slow-movers!” – even though it’s corroborated by MAE and MRPS, “We must improve Hobbies 2! Slow-movers are performing worse than fast-movers!” is backed by WMAPE and NMRPS. The mere inconsistency of these conclusions is already striking and shows that the judgement of forecast quality across departments is not straightforward. The reason for the reversed conclusions is, again, the precision scaling of metrics: When increasing the selling rate, the typical values that unnormalized metrics (MAE, MRPS) achieve increase, the typical values of normalized (WMAPE, NMRPS) metrics decrease.
@@
A direct interpretation of the forecast metrics across departments (Bower, 2023) (or any other way of segregating the data) is deeply problematic and is doomed to misalign efforts for forecast improvement. Since the “effect” that one sees is infected by precision scaling, we propose to designate such distortion as the naïve scaling trap.
@@ 3.1 Revealing and avoiding the naïve scaling trap
The reason why the metrics in Table 1 amount to different values in the different departments lies in the scaling of the variance of the Poisson distribution with the rate $\mu$: Each individual prediction $r_{j}$ comes with a different expectation of how $s_{j}$ should be distributed (Eq. (4)), and which metric value $M(r_{j},s_{j})$ should be achieved on average. The widespread assumption that the same metric value is achievable for each category is therefore flawed. The value of any metric primarily reflects the unavoidable noise that also the sharpest-possible forecast exhibits. Only secondarily does the metric value reflect forecast quality.
It is therefore useful to compute the expected value of the metric, given the predictions, under the assumed sharpest possible forecast. These ideal values constitute a context for the actually achieved metrics. A marginal discrepancy would confirm that the forecast is (almost) as good as it could possibly be. A large difference motivates an investigation whether it is possible to close that gap.
If all predictions are aggregated globally, the resulting expectation value of the metric will be mostly borne by fast-movers. Therefore, a segregation by selling rate is useful: We group the pairs of predictions $r_{j}$ and observations $s_{j}$ into buckets, the bucket of a pair is chosen according to the prediction value $r_{j}$. If one chose the outcome $s_{j}$ to define the bucket, one would suffer from a hindsight selection bias (Tichy and Feindt, 2022), which leads to the forecaster’s dilemma (Lerch et al., 2017; Tichy, 2023b, 2024). Logarithmically spaced buckets are useful in retail to avoid cases in which one bucket contains the majority of all predictions, i.e., we group by
@@
where we round to integers such that there are $n_{\text{bins}}$ bins between two powers of 10. For example, for $n_{\text{bins}}=4$, the $R_{j}$ that can take the values $0,0.25,0.5,0.75,1,1.25$ etc. Buckets are referred to by their common rounded logarithmic prediction value $R$. In practice, we have found that $n_{\text{bins}}=2$ constitutes the absolute minimum of buckets, and choosing $n_{\text{bins}}=5$ is typically a reasonable choice. The computed ratings also weakly depend on the number of buckets, such that it should be chosen as a constant within one analysis.
The predictions $r_{j}$ that belong to the same bucket, indexed by $R$, come with similar expectation values of any metric. Ideally, when the prediction is truly Poissonian, we expect the achieved average metric in a bucket to match the metric’s computed expectation value. The number of bins $n_{\text{bins}}$ should thus solve the compromise between having sufficiently many predictions and observations in one bucket so we can expect measurement and expectation value to match approximately (i.e. a possible deviation to be considered significant), while ensuring that all predictions in the bucket behave approximately equally. Figure 2 shows the result of this procedure for the baseline model as circles and $n_{\text{bins}}=4$, and superimposes the expectation value of the metrics under the Poisson distribution as solid line. The circle sizes reflect the square-root of the number of measurements in order to make small buckets still visible. The positions of the circles match the solid lines, confirming the Poissonian nature of our baseline forecast from Figure 1. Figure 3 differentiates further by retail department for the NMRPS: The differences in the metric values exhibited in Table 1 are driven by the differences in the distribution of rates, not by a different performance. For a given rate bucket $R$, the performance is the same for all departments, but the different buckets have different populations. Skepticism that calls into question the equal model performance in the different departments, which might have been induced by the different metric values in Table 1, can now be clearly dismissed.
Figure 2: Ideal Poisson metric values (solid lines) and achieved metric values for baseline model (circles). The size of the circles reflects the square root of the number of sales grouped in one bucket.
@@
The intricate shapes that metric values exhibit as a function of the predicted rate in Figure 2 make comparisons of traditional metrics between strata of data (Bower, 2023) deeply problematic. The “best case” value of a metric depends strongly on the prediction itself. Faster-selling departments naturally achieve higher values of MAE and MRPS, and lower values of WMAPE and NMRPS, as exemplified by Figure 3. Therefore, statements like “we achieve 38% WMAPE on Foods 3, but 78% WMAPE on Hobbies 2, we need to improve Hobbies 2” are almost always infected by the naïve scaling trap and need to be dismissed.
3.2 Necessity for imperfect reference forecasts

### v5

@@ 2.2 Corrections to the Poisson distribution
In practice, the Poisson assumption laid out above is never perfectly fulfilled. Several effects both increase and decrease the width of the ideally achievable distribution. In the first place, perfectly knowing the number of visitors $N$ and the average probability $p$ to buy an item is a strong assumption – not all factors that influence demand can be known. Hence, this item-level buy rate $\mu=N\cdot p$ is itself randomly distributed. If that distribution is gamma, the demand distribution becomes gamma-Poisson (negative-binomial).
The forecast could exploit personal information about customers. Instead of an average probability $p$ for the individual buy, this might be personalized to $p_{1},\dots,p_{N}$. Le Cam’s theorem (Le Cam, 1960) sets a bound to the total variation distance between the resulting Poisson-binomial and Poisson distributions. Under plausible assumptions, the potential of added forecast sharpness via personalization is severely limited: An extraordinary degree of individualization (individual customer buy probabilities need to convey a lot of information, i.e. $p_{j}\ll 1$ or $p_{k}\approx 1$) is necessary to achieve a clearly-sharper-than-Poisson distribution.
The assumption that only one item is bought per customer is also slightly broken in practice: At Sainsbury’s UK, about 87.5% of sold items are bought alone, 9.8% are bought in pairs of two pieces, about 1.5% in triplets of three pieces, four or more pieces are also occurring. The variance of the resulting distribution is thereby increased. In grocery retail, however, the effect of multi-buys amounts to a manageable correction.
Finite stocks censor demand values that are larger than the available stocks. This right censoring effectively decreases the width of the observed distribution.
A priori, it is difficult to estimate the strength of these effects. For the datasets that we employed – M5 and Sainsbury’s – we will see, ex post, that the Poisson limit is a first-order benchmark, and that the effects mentioned in this section can be considered “second-order terms”. If a more suitable benchmark distribution is available, which may include effects such as over-dispersion, zero-inflation, and right censoring, our method can easily be adapted, mutatis mutandis.
To focus on our main contribution, the rating of forecasts across selling rates, our working assumption is that a Poisson forecast constitutes the sharpest possible prediction.
@@ 2.3 Forecasting model
Under the Poisson assumption, a product-location-day combination $j$ ($1\leq j\leq n$) is governed by a true selling rate $t_{j}$, which will remain unknown and is only under control in the numerical experiments here. The observed sales $s_{j}$ are samples from the Poisson distribution with rate $\mu=t_{j}$:
@@
The forecast for that product-location-day $j$ is a selling rate $r_{j}$. The forecaster is self-confident and believes that the forecasted rates are perfectly known Poisson-rates, such that each $s_{j}$ behaves as if it were drawn from that Poisson distribution. Our considerations are independent of how the forecaster constructed the $r_{j}$: They might use a time-series approach, a machine-learning model, hybrid approaches, or any other method (Petropoulos, F. et al., 2022).
@@ 2.4 Evaluation metrics
To evaluate the quality of their forecast, the forecaster relates the observed sales $\vec{s}=(s_{1},\dots,s_{n})$ to their previously predicted rates $\vec{r}=(r_{1},\dots,r_{n})$ by evaluating metrics $M(\vec{r},\vec{s})$.
@@ 2.4.1 Loss functions on point estimates
A popular metric is Mean Absolute Error (MAE),
@@
where the optimal point estimator for MAE, the median of the predicted distribution, replaces the mean $r_{j}$ (Schwertman et al., 1990; Gneiting, 2011). That is, the forecaster is aware that their forecast $r_{j}$ is not a universal best guess suitable for every metric, but that a metric-dependent point estimate is required (Kolassa, 2020). MAE naturally takes larger values for larger predictions, a first attempt towards a scale-free metric is the dimensionless normalized MAE (Kolassa and Schütz, 2007), often called Weighted Mean Absolute Percentage Error (WMAPE),
@@
The still-popular Mean Absolute Percentage Error (MAPE) comes with many flaws (Hyndman and Koehler, 2006; Kolassa and Martin, 2011; Kolassa and Schütz, 2007). In particular, it is undefined for observations $s=0$, whose treatment heavily affects the achieved MAPE-values (Kolassa, 2023a), complemented by the particularly complicated precision scaling of MAPE (Tichy, 2023a).
@@ 2.4.2 Ranked Probability Score
MAE has an immediate business interpretation: A planner who orders $\text{median}(r_{j})$ pieces will, on average, have an excess or under-stock of MAE pieces. Using the median of the distribution, however, leads to unpleasant discontinuous behavior (the median jumps from 0 to 1 at the rate $r=\log 2\approx 0.693$). To provide a business interpretation while acknowledging the probabilistic nature of the forecast, the forecaster also evaluates the Mean Ranked Probability Score (MRPS), which is the mean over all product-location-days of the discrete Ranked Probability Score (RPS) (Epstein, 1969; Gneiting and Raftery, 2007; Tichy, 2022).
@@
MRPS generalizes MAE from simple point estimates to distributions: For a deterministic zero-width forecast that states that $r_{j}$ occurs with certainty, the first term in Eq. (6) becomes $|r_{j}-s_{j}|$} and second term vanishes (Gneiting and Raftery, 2007). Our main working horse in this article, the Normalized Mean Ranked Probability Score (NMRPS), is then defined as
@@
Precision scaling is ubiquitous, it affects (N)MRPS and all other metrics in a similar way as MAE/WMAPE.
@@ 2.5 Baseline model and naïve scaling trap
As test-bed of our method, we use the dataset for the validation period, 2016-04-25 to 2016-05-22, of the M5-competition (Makridakis et al., 2022b, a). To obtain a maximally sharp and calibrated placeholder for a perfect baseline forecast, we construct predictions $r^{\text{baseline}}_{j}$ via in-sample Expectation Maximization (EM) of the Poisson likelihood, applied on the entire set of sales $\vec{s}$ (see Supplemental Material). In short, the Poisson-rates $r_{j}$ that are generated by this procedure could plausibly have have produced the outcomes $s_{j}$.
A forecaster is interested in the overall performance of their model (“is that a good model?”), and in systematic patterns in the performance (“are we doing better in Hobbies or in Household?”), and therefore evaluates the forecast not only globally, but also segregated by department (or by other properties known at the moment of the forecast).
@@
In Table 1, for the total, and for each of three selected departments (“Household 1”, “Hobbies 2”, “Foods 3”), all achieved metrics are finite for the baseline forecast. Remember that the forecast was fabricated to be ideally Poissonian for the entire assortment – by construction, it has the same quality in every subset of data. The result of a correct analysis should produce the result “there is no significant difference in forecast quality between the departments”. Hence, the following intuitive conclusions are erroneous: “We need to improve Foods 3! Fast-movers are performing worse than slow-movers!” – even though it’s corroborated by MAE and MRPS, “We must improve Hobbies 2! Slow-movers are performing worse than fast-movers!” is backed by WMAPE and NMRPS. The striking inconsistency of these conclusions shows that the judgement of forecast quality across departments is not straightforward.
@@
A direct interpretation of the forecast metrics across any segregation of the data (Bower, 2023) is deeply problematic and is doomed to misalign efforts for forecast improvement. Since the “effect” that one sees is infected by precision scaling, we propose to designate such distortion as the naïve scaling trap.
@@ 3.1 Revealing and avoiding the naïve scaling trap
The reason why the metrics in Table 1 amount to different values in the different departments lies in the scaling of the variance of the Poisson distribution with the rate $\mu$: Each individual prediction $r_{j}$ comes with a different expectation of how $s_{j}$ should be distributed (Eq. (3)), and which metric value $M(r_{j},s_{j})$ should be achieved on average. The widespread assumption that the same metric value is achievable for each category is therefore flawed.
It is useful to compute the expected value of the metric, given the predictions, under the assumed sharpest possible forecast. These ideal values constitute a baseline comparison for the actually achieved metrics. A marginal discrepancy would confirm that the forecast is almost as good as possible. A large difference motivates an investigation whether it is possible to close that gap.
If all predictions are aggregated globally, the resulting expectation value of the metric will be mostly borne by fast-movers. We therefore segregate by selling rate: We group the pairs of predictions $r_{j}$ and observations $s_{j}$ into buckets, the bucket of a pair is chosen according to the prediction value $r_{j}$. Choosing the outcome $s_{j}$ as identifier would induce the hindsight selection bias (Tichy and Feindt, 2022; Lerch et al., 2017; Tichy, 2023b, 2024a). Logarithmically spaced buckets avoid that one bucket contains the majority of all predictions, i.e., we group by
@@
where we round to integers such that there are $n_{\text{bins}}$ bins between two powers of 10. For example, for $n_{\text{bins}}=4$, the $R_{j}$ that can take the values $0,0.25,0.5,0.75,1,1.25$ etc. Buckets are referred to by their common rounded logarithmic prediction value $R$. The size and number of the buckets should then ensure a good balance between precision (many small, homogenous buckets) and sufficient statistics per bucket (few large more inhomogeneous buckets); in practice, $n_{\text{bins}}\approx 5$ is a reasonable choice.
The predictions $r_{j}$ that belong to the same bucket, indexed by $R$, come with similar expectation values, for any metric. Ideally, when the prediction is truly Poissonian, we expect the achieved average metric in a bucket to match the metric’s computed expectation value.
Figure 1 illustrates the result of this procedure for the baseline model as circles and $n_{\text{bins}}=4$, and superimposes the expectation value of the NMRPS under the Poisson distribution as solid lines. The positions of the circles match the lines, confirming the Poisson performance of our baseline model. Figure 2 differentiates further by retail department: The differences in the metric values exhibited in Table 1 are driven by the differences in the distribution of rates among the departments, not by a different performance. For a given rate bucket $R$, the performance is the same for all departments, but the different buckets have different populations, which leads to a different overall metric value. Skepticism that calls into question the equal model performance in the different departments, which might have been induced by the different metric values in Table 1, can now be rejected.
Figure 1: Ideal Poisson NMRPS (solid lines) and achieved NMRPS for baseline model (circles). The size of the circles reflects the square root of the number of sales grouped in one bucket.
@@
The intricate shapes that the NMRPS exhibits as a function of the predicted rate in Figure 1 makes comparisons of traditional metrics between strata of data (Bower, 2023) deeply problematic. The “best case” value of a metric depends strongly on the prediction itself. Faster-selling departments naturally achieve higher values of MAE and MRPS, and lower values of WMAPE and NMRPS, as exemplified by Figure 2. Therefore, statements like “we achieve 38% WMAPE on Foods 3, but 78% WMAPE on Hobbies 2, we need to improve Hobbies 2” are infected by the naïve scaling trap and need to be dismissed.
3.2 Imperfect reference forecasts